# A unified linear intrinsic alignment model for elliptical and disc galaxies and the resulting ellipticity spectra

### v1

@@ Acknowledgements
Fig. 7 shows the effect of an isotropic tidal field $\Delta\Phi$ to the orbit of a star at $r=r_{s}$. As expected, there is a change in size of the galactic disc, analogous to the effect on elliptical galaxies (Ghosh et al.(2021)Ghosh, Durrer, and Schaefer) or, equivalently, to the effect of lensing convergence (Bartelmann and Schneider(2001)). In both cases, galaxies would appear larger in low-density environments and smaller in high-density environments. Visually, the effect of an isotropic tidal field on the size of the disc seems to be smaller compared to the shear depicted in Fig. 1, but one perceives only half of the effect as both semi-axes are changed by the same factor $21+\Omega^{2}/\omega^{2}/$ 2 and unlike in the case of ellipticity, $21+\Omega^{2}/\omega^{2}/$ 2 relative to $21-\Omega^{2}/\omega^{2}/$ 2. The same argument as for disc-averaged ellipticities would apply here as well, as less tightly bound stars at larger radii would experience a stronger tidal enlargement of their orbital radii. In consequence, one should expect an size alignment effect with an equal amplitude as the ellipticity alignment effect, in analogy to the result for size alignments in elliptical galaxies (Ghosh et al.(2021)Ghosh, Durrer, and Schaefer). As the size-alignments for elliptical galaxies are difficult to measure even for surveys such as Euclid’s, it is doubtful if the analogous measurement for disc galaxies could yield a statistically significant detection, extrapolating the results by Ghosh et al.(2021)Ghosh, Durrer, and Schaefer. Repeating the estimates in Sect. 3 in conjunction with the results of Ghosh et al.(2021)Ghosh, Durrer, and Schaefer yields the result that correlations between size perturbations and weak lensing convergence are likewise difficult to measure with data from a Euclid-like survey.

### v2

@@ Acknowledgements
Fig. 7 shows the effect of an isotropic tidal field $\Delta\Phi$ to the orbit of a star at $r=r_{s}$. As expected, there is a change in size of the galactic disc, analogous to the effect on elliptical galaxies (Ghosh et al., 2021) or, equivalently, to the effect of lensing convergence (Bartelmann and Schneider, 2001). In both cases, galaxies would appear larger in low-density environments and smaller in high-density environments. Visually, the effect of an isotropic tidal field on the size of the disc seems to be smaller compared to the shear depicted in Fig. 1, but one perceives only half of the effect as both semi-axes are changed by the same factor $21+\Omega^{2}/\omega^{2}/$ 2 and unlike in the case of ellipticity, $21+\Omega^{2}/\omega^{2}/$ 2 relative to $21-\Omega^{2}/\omega^{2}/$ 2. The same argument as for disc-averaged ellipticities would apply here as well, as less tightly bound stars at larger radii would experience a stronger tidal enlargement of their orbital radii. In consequence, one should expect a size alignment effect with an equal amplitude as the ellipticity alignment effect, in analogy to the result for size alignments in elliptical galaxies (Ghosh et al., 2021). Correlations between size and alignments have also been highlighted previously in literature (Vlah et al., 2020; Ciarlariello et al., 2015; Ciarlariello and Crittenden, 2016; Johnston et al., 2023). As the size-alignments for elliptical galaxies are difficult to measure even for surveys such as Euclid’s, it is doubtful if the analogous measurement for disc galaxies could yield a statistically significant detection, extrapolating the results by Ghosh et al. (2021). Repeating the estimates in Sect. 3 in conjunction with the results of Ghosh et al. (2021) yields the result that correlations between size perturbations and weak lensing convergence are likewise difficult to measure with data from an Euclid-like survey.